If an economy is growing at 4% a year, how long will it take for its income to double?
Answer #1

Growth rate = 4% = 0.04

Let.

Present Income = Y

So Double Income = 2Y

Time for it to double = n years

From Growth formula we know

Future Income = Present Income 1 + Growth rate)

So,

2Y = Y(1 + 0.04)n

2 = (1.04)n

Taking log on both sides :

log(2) = n * log(1.04)

n = log(2)/log(1.04) = 0.301/0.017

n = 17.7 years ~ 18 years

So, It will take around 18 years for the economy to double its income at growth of 4% per year.